# Richelieu: Self-Evolving LLM-Based Agents for AI Diplomacy

### v1

@@ 4.3 Negotiator and Actor
By chatting with other players, the goal of the negotiation is to update the social belief according to the received words and reach the sub-goal by manipulating other’s intentions, such as securing cooperative agreements with other nations, terminating ongoing conflicts with a specific country, or deterring the formation of alliances directed against its interests.[37, 67] However, it is difficult to reach a consensus, as the interests and strategies of the various nations often conflict, and trust between players can be scarce, making it challenging to establish and maintain cooperative agreements. In this case, we argue that the negotiator should identify the true intentions and relationship of the opponent before generating the words for the negotiation.
To fully utilize the power of LLMs, we construct a social reasoning flow for negotiation, as shown in Figure 3. During the negotiation process, we guide Richelieu to consider the veracity of what other players said and their true intentions, and in conjunction with our established sub-goals and analysis of our relationships with other players, to negotiate and form alliances with potential allies and attempt to deceive enemies.[60, 35]
@@ 4.4 Memory Management and Self-Evolution
This memory is the foundation of the framework that accumulates the historical experience of the agent and summarizes them for other modules.[17, 30, 66, 23] It supports other modules, such as planner and negotiator, to provide long-tail experiences.
@@
Acquisition Experience via Self-Play Games. Self-play allows the agent to accumulate more experiences for self-evolution.[33, 69] After training, when Richelieu is faced with a certain state, it can draw on a larger pool of similar historical experiences. Diverse evaluations enable Richelieu to reflect more comprehensively on the strategies it currently devises, leading to a stronger optimization of decision making. As self-play continues, the acquisition of new and better historical experiences by Richelieu will diminish. This means that Richelieu’s capabilities will not improve indefinitely. At the same time, as the memory grows, selecting appropriate historical experiences becomes a new challenge. The chosen m experience $\vec{\eta_{t}}$ may be almost identical, which could actually reduce the amount of useful information available to Richelieu. As shown in Figure 5, Richelieu’s performance against Cicero [6] becomes better with increasing training iterations. With the accumulation of experiences, Richelieu’s win rate exhibited a steady increase with accumulated training iterations, ultimately plateauing at a stable performance level. In contrast, the defeated rate showed a consistent decrease, approaching an asymptotic value. These observations confirm the effectiveness of self-play in Richelieu’s evolution.
@@ 5 Experiment
In the experiments, our goal is to answer the following questions: 1) Mastery of Non-Press Diplomacy: Can our agent master the non-press diplomacy against baselines? 2) Competing with State-of-the-Art: Can our agent surpass the performance of the current state-of-the-art agents in press diplomacy? 3) Compatibility with LLMs: Can our self-evolving framework be compatible with different LLMs? 4) Contribution of Framework Modules: Do the individual modules within our framework contribute to the overall improvement of our agent’s performance?
@@ 5.1 Experimental Setup
Environment. The widely-used open source Diplomacy game platform introduced by [39] is adopted for evaluating Richelieu against other models. It is easy to switch between no-press (with communication/negotiation between players) and press (no communication between players) games based on this platform, facilitating comparison on both settings. The platform also contains over 10,000 human game data on which previous approaches are trained. Note that our method does not need them. In each game, a model will play the role of one randomly selected country to compete against countries controlled by other methods. It wins if occupying all the supply centers and loses vice versa.
Evaluation Metrics. We evaluate the models based on the results of multiple rounds of games. In each round, the model is randomly assigned a country to control. Typically, 1000 rounds are played to obtain the average results. We evaluate the models in two metrics. One is based on the win rate, Most SC rate, survived rate, and defeated rate. There are four possible outcomes for each country in the game. If a country loses all its supply centers (SC), it is eliminated and recorded as “defeated". If a country occupies 18 or more out of 34 supply centers, the game ends, and that country is recorded as “win", while other countries are recorded as “defeated". In other cases, the game ends in a draw. The country with the most supply centers is recorded as “Most SC", the countries that have been eliminated are recorded as “defeated", and the other countries are recorded as “Survived". The other is based on the scores obtained by the models after multiple rounds of competition. To compare the capabilities of multiple models, we use C-Diplo Argir[4], a scoring system. This system is used in many international diplomacy competitions. The scoring method is as follows: If a player wins by occupying 18 or more supply centers, the player scores 93 points, and each of the other six players scores 1 point. If the game ends in a draw, the player with the most centers scores 37 points. The second player with the most centers scores 14 points. The third player with the most centers scores 7 points. Each player scores 1 point per center owned. Each player also scores 1 point for participating. In this way, regardless of the game outcome, a total of 99 points will be distributed among the players in each game.
Baselines. We select six previous models as baselines for comparison. Among them, Cicero[6] by Meta is a diplomacy model with a negotiation module. The other five are no-press diplomacy models, including the SL-DipNet and RL-DipNet [39], the BRPI [3], the SearchBot [18], and the DORA[5].
5.2 Results
Massively Play with Baselines on no-press setting. We let Richelieu compete with the other six models including Cicero[6], SL-DipNet and RL-DipNet [39], BRPI [3], SearchBot [18], and DORA[5] on No-Press Diplomacy, in which players make moves without communication. Figure  4 indicates that Richelieu outperforms other previous models relying on human game data. In contrast, Richelieu does not need such data but outperforms these methods by a clear margin, which demonstrates the outstanding planning capability of Richelieu.
@@ 5.2 Results
Generalization of self-evolving framework to diverse LLMs. Llamma 3 To demonstrate the effectiveness of our framework in a variety of LLM, we conducted experiments using four models: GPT4.0, ERNIE Bot, Spark Desk, and Llamma 3. The experimental results show that, despite variations in Richelieu’s performance due to the inherent differences in the capabilities of these LLMs, as illustrated in Figure 5, our framework and training approach significantly enhance the capabilities of all large language models. After training, the win rate using GPT4.0 increased from 1.5% lower than Cicero’s to about 0.7% higher than Cicero’s. The win rate using llama3 increased from 2.3% lower than Cicero’s to almost equal to Cicero’s. The win rates using Models Spark Desk and ERNIE Bot increased from 3% and 4% lower than Cicero’s to 0.7% and 1.6% lower than Cicero’s, respectively. This indicates the generalization of a self-evolving framework to various LLMs.
@@
Ablation Study. We conduct comprehensive ablation studies on Richelieu by analyzing the benefit of incorporating Richelieu’s various modules, like planners or memory, into basic LLMs. The results are shown in Table 2. As can be seen, direct use of vanilla LLM yields relatively poor results. Richelieu’s performance obtains steady and significant improvement by incorporating each individual module. This indicates that Richelieu is able to leverage other players’ actions during decision-making and consider both short-term and long-term benefits. Additionally, Richelieu’s negotiation ability has been significantly improved, allowing it to effectively express intentions to cooperate with other players and avoid deception during negotiations. And after self-play, Richelieu’s experience makes it perform better.
@@
6 Example Cases
Figure 6: An example case of negotiating with a nation to form an alliance to confront a strong enemy.
As is shown in Figure  6, Richelieu controls France. At this time, France is at war with Austria over control of the Apennine Peninsula. However, Russia is on the verge of victory in its war against Turkey, which will lead to significant territorial expansion for Russia. Although France and Russia currently do not share a border, are not at war, and have no conflicts of interest, Richelieu foresees Russia becoming the most threatening enemy in the future. Therefore, Richelieu sets a sub-goal of weakening Russia.
In the subsequent negotiation phase, Richelieu proactively proposes ending the war with Austria, despite holding an advantage in this conflict. Richelieu promises Austria that if it ceases hostilities and attacks Russia, Richelieu will assist Austria in defending against any attacks from England. The negotiations are successful. Austria accepted Richelieu’s proposal, and the two countries reached an agreement to exchange the supply centers of Napoli and Munich.
During the action phase, Austria moves its troops from Venice to Apulia in preparation for capturing Napoli in the next turn, while the rest of its forces are repositioned to the eastern regions bordering Russia to defend against Russian attacks and compete for supply centers. French units occupy Munich and prepare to advance on Russian territories such as Berlin. Meanwhile, French units support Austria in the Holland and Belgium regions.
@@ 8 Limitations and Future Work
The potential applications of such an AI agent are vast, ranging from simulated diplomatic environments to real-world assistance and analysis. In future research, we intend to develop a more realistic game space, characterized by incomplete information and multi-player games, to enhance and refine our model further. We will also extend the framework to other multi-agent scenarios, including embodied interactions [75, 11, 9], sensor networks [54, 61, 38, 31], and video games [49, 34]. This framework can also be employed to develop various applications. For instance, in the fields of economics and finance, we intend to utilize it to create business analytics and negotiation models.
@@ A.1 Rules of Diplomacy Game
  • The number of units a country can have cannot exceed the number of supply centers it controls. If the number of supply centers decreases, excess units must be disbanded. Each autumn, new units can be built at supply centers. Coastal supply centers can produce fleets or armies, while others can only produce armies. [22]
@@ A.2 Domain Knowledge
Richelieu can adopt a strategy of allying with distant countries while attacking neighboring ones to occupy adjacent territories and achieve rapid expansion. Richelieu should pay attention to the Balance of Power by forming alliances with other countries or supporting weaker states to prevent any single country or alliance from becoming too powerful. [12] To this end, Richelieu can also adopt a strategy of attacking distant countries while allying with nearby ones, sacrificing short-term benefits to avoid the emergence of future hegemonic states that could threaten his own survival. When facing multiple enemies, Richelieu can find ways to divide other countries and incite wars among them. Whether in offense or defense, Richelieu should actively choose suitable allies. Richelieu can also introduce a third party to achieve goals such as ceasefire, alliance, or joint attack. To achieve alliances or ceasefires, Richelieu can sacrifice some interests to the other party as long as the ultimate benefits are greater. Others may lie and deceive [27]; their words in negotiations are not binding. Richelieu must avoid being deceived or betrayed. At the same time, Richelieu can also actively deceive others to achieve his own goals.[42, 2]

### v2

@@ 4.3 Negotiator and Actor
By chatting with other players, the goal of the negotiation is to update the social belief according to the received words and reach the sub-goal by manipulating other’s intentions, such as securing cooperative agreements with other nations, terminating ongoing conflicts with a specific country, or deterring the formation of alliances directed against its interests.Noh and Chang (2024); Zhan et al. (2024) However, it is difficult to reach a consensus, as the interests and strategies of the various nations often conflict, and trust between players can be scarce, making it challenging to establish and maintain cooperative agreements. In this case, we argue that the negotiator should identify the true intentions and relationship of the opponent before generating the words for the negotiation.
To fully utilize the power of LLMs, we construct a social reasoning flow for negotiation, as shown in Figure 3. During the negotiation process, we guide Richelieu to consider the veracity of what other players said and their true intentions, and in conjunction with our established sub-goals and analysis of our relationships with other players, to negotiate and form alliances with potential allies and attempt to deceive enemies.Xia et al. (2024); Moghimifar et al. (2024)
@@ 4.4 Memory Management and Self-Evolution
This memory is the foundation of the framework that accumulates the historical experience of the agent and summarizes them for other modules.Gao and Zhang (2024); Li et al. (2024a); Yu et al. (2024); Hou et al. (2024) It supports other modules, such as planner and negotiator, to provide long-tail experiences.
@@
Acquisition Experience via Self-Play Games. Self-play allows the agent to accumulate more experiences for self-evolution.Liu et al. (2024); Zhang et al. (2024a) After training, when Richelieu is faced with a certain state, it can draw on a larger pool of similar historical experiences. Diverse evaluations enable Richelieu to reflect more comprehensively on the strategies it currently devises, leading to a stronger optimization of decision making. As self-play continues, the acquisition of new and better historical experiences by Richelieu will diminish. This means that Richelieu’s capabilities will not improve indefinitely. At the same time, as the memory grows, selecting appropriate historical experiences becomes a new challenge. The chosen m experience $\vec{\eta_{t}}$ may be almost identical, which could actually reduce the amount of useful information available to Richelieu. As shown in Figure 5, Richelieu’s performance against Cicero (Bakhtin et al., 2022) becomes better with increasing training iterations. With the accumulation of experiences, Richelieu’s win rate exhibited a steady increase with accumulated training iterations, ultimately plateauing at a stable performance level. In contrast, the defeated rate showed a consistent decrease, approaching an asymptotic value. These observations confirm the effectiveness of self-play in Richelieu’s evolution.
@@ 5 Experiment
In the experiments, our goal is to answer the following questions: 1) Mastery of Non-Press Diplomacy: Can our agent master the non-press diplomacy against baselines? 2) Competing with State-of-the-Art: Can our agent surpass the performance of the current state-of-the-art agents in press diplomacy? 3) Compatibility with LLMs: Can our self-evolving framework be compatible with different LLMs? 4) Contribution of Framework Modules: Do the individual modules within our framework contribute to the overall improvement of our agent’s performance? 5) Social Reasoning: Can Richelieu accurately infer the true intentions of other players and reasonably determine the relationships of ally or enemy with them?
@@ 5.1 Experimental Setup
Environment. The widely-used open source Diplomacy game platform introduced by (Paquette et al., 2019) is adopted for evaluating Richelieu against other models. It is easy to switch between no-press (with communication/negotiation between players) and press (no communication between players) games based on this platform, facilitating comparison on both settings. The platform also contains over 10,000 human game data on which previous approaches are trained. Note that our method does not need them. In each game, a model will play the role of one randomly selected country to compete against countries controlled by other methods. It wins if occupying all the supply centers and loses vice versa.
@@
Evaluation Metrics. We evaluate the models based on the results of multiple rounds of games. In each round, the model is randomly assigned a country to control. Typically, 1000 rounds are played to obtain the average results. We evaluate the models in two metrics. One is based on the win rate, Most SC rate, survived rate, and defeated rate. There are four possible outcomes for each country in the game. If a country loses all its supply centers (SC), it is eliminated and recorded as “defeated". If a country occupies 18 or more out of 34 supply centers, the game ends, and that country is recorded as “win", while other countries are recorded as “defeated". In other cases, the game ends in a draw. The country with the most supply centers is recorded as “Most SC", the countries that have been eliminated are recorded as “defeated", and the other countries are recorded as “Survived". The other is based on the scores obtained by the models after multiple rounds of competition. To compare the capabilities of multiple models, we use C-Diplo Argir(Archer, 2024), a scoring system. This system is used in many international diplomacy competitions. The scoring method is as follows: If a player wins by occupying 18 or more supply centers, the player scores 93 points, and each of the other six players scores 1 point. If the game ends in a draw, the player with the most centers scores 37 points. The second player with the most centers scores 14 points. The third player with the most centers scores 7 points. Each player scores 1 point per center owned. Each player also scores 1 point for participating. In this way, regardless of the game outcome, a total of 99 points will be distributed among the players in each game.
Baselines. We select six previous models as baselines for comparison. Among them, Cicero(Bakhtin et al., 2022) by Meta is a diplomacy model with a negotiation module. The other five are no-press diplomacy models, including the SL-DipNet and RL-DipNet (Paquette et al., 2019), the BRPI (Anthony et al., 2020), the SearchBot (Gray et al., 2020), and the DORA(Bakhtin et al., 2021). We also build a LLM-based agent, AutoGPT Yang et al. (2023a). In experiments, we set a temperature of 0.3 to ensure a relatively stable generation of LLM policies. The overall reasoning framework also ensure the stability and consistency in the AI agent’s performance.
5.2 Results
Massively Play with Baselines on no-press setting. We let Richelieu compete with the other six models including Cicero(Bakhtin et al., 2022), SL-DipNet and RL-DipNet (Paquette et al., 2019), BRPI (Anthony et al., 2020), SearchBot (Gray et al., 2020), and DORA(Bakhtin et al., 2021) on No-Press Diplomacy, in which players make moves without communication. Figure  4 indicates that Richelieu outperforms other previous models relying on human game data. In contrast, Richelieu does not need such data but outperforms these methods by a clear margin, which demonstrates the outstanding planning capability of Richelieu.
@@ 5.2 Results
Although Richelieu’s win rate improvement compared to Cicero is not significant, the relative value of the improvement is quite large. Moreover, the main reason for the modest improvement is that in the seven countries, there are three or four controled by Richelieu with similar abilities, which often results in the game ending in a draw. Moreover, we observed a large gap by comparing the scores the agents gained in the massively play with baselines on no-press setting showed in Figure 4. Our agent’s score is about 10% higher than Cicero’s.
Table 2: The results of our method playing against AutoGPT.
[…]
Richelieu_1 9.30% 18.20% 37.90% 34.60%
Richelieu_2 9.90% 19.40% 37.70% 33.00%
Richelieu_3 8.10% 17.40% 39.20% 35.30%
AutoGPT_1 1.20% 4.60% 32.40% 61.80%
AutoGPT_2 1.20% 4.20% 34.40% 60.20%
AutoGPT_3 1.50% 4.00% 32.50% 62.00%
AutoGPT_4 2.60% 3.60% 32.30% 61.50%
Richelieu 9.10% 18.33% 38.27% 34.30%
AutoGPT 1.63% 4.10% 32.90% 61.37%
Play against AutoGPT on press setting. We further built an LLM-based agent using AutoGPT and compared it with our agent. In the testing, we randomly select three countries to be controlled by Richelieu, and the other four countries to be controlled by AutoGPT. Note that the agent controls each country independently. The results are showed in table  2. The results show that our model outperforms the existing LLM baseline.
Generalization of self-evolving framework to diverse LLMs.
Llamma 3 To demonstrate the effectiveness of our framework in a variety of LLM, we conducted experiments using four models: GPT4.0, ERNIE Bot, Spark Desk, and Llamma 3. As the number of training iterations increases, Richelieu’s win rate steadily improves while the defeated rate declines, ultimately reaching a relatively stable outcome. This suggests that our self-play method is effective. After training, the win rate using GPT4.0 increased from 1.5% lower than Cicero’s to about 0.7% higher than Cicero’s. The win rate using llama3 increased from 2.3% lower than Cicero’s to almost equal to Cicero’s. The win rates using Models Spark Desk and ERNIE Bot increased from 3% and 4% lower than Cicero’s to 0.7% and 1.6% lower than Cicero’s, respectively. The experimental results show that, despite variations in Richelieu’s performance due to the inherent differences in the capabilities of these LLMs, as illustrated in Figure 5, our framework and training approach significantly enhance the capabilities of all large language models.This indicates the generalization of a self-evolving framework to various LLMs.
In order to demonstrate the effect of the memory from the self-play game on the strategy of our agent, we found two turns with similar states in different rounds, one before self-play and the other after. The cases are showed in Appendix B.1.
@@
Ablation Study. We conduct comprehensive ablation studies on Richelieu by analyzing the benefit of incorporating Richelieu’s various modules, like planners or memory, into basic LLMs. The results are shown in Table 3. As illustrated in Figure  5, while the enhanced alignment in LLMs indeed boosts performance (GPT-4.0 is better than others), we observed that a vanilla GPT-4.0 still falls short in AI diplomacy without our framework, as can be seen in Table  3. Richelieu’s performance obtains steady and significant improvement by incorporating each individual module. This indicates that Richelieu is able to leverage other players’ actions during decision-making and consider both short-term and long-term benefits. Additionally, Richelieu’s negotiation ability has been significantly improved, allowing it to effectively express intentions to cooperate with other players and avoid deception during negotiations. And after self-play, Richelieu’s experience makes it perform better. These indicate that the alignment in LLMs lays a foundation, but our approach is key to unlocking the models’ potential in social simulation.
Table 4: The success rate to identify the social relationship and infer others’ intentions.
GPT-4.0 Llama3
relationship 85.74% 85.52%
intention(sub-goal) 74.67% 74.11%
Social Reasoning. We conduct an experiment to evaluate the success rate that the agent can successfully identify the social relationship and infer others’ intentions. As the baselines do not explicitly model the relationship and intention, we can not directly access the ground truth for evaluation. Instead, we let all players use our agent but with different LLMs, i.e., 4 countries use GPT-4.0 and 3 countries use Llama3. The accuracy is reported in table  4. We can see that the accuracy of social reasoning is consistent with the overall performance of the agent, indicating the effectiveness of social reasoning.
@@ 7 Limitations and Future Work
Our framework is capable of applying to most social interaction tasks. Most components in our framework can be easily generalized to a new task by modifying the content. Social reasoning enables the agent to handle complex and dynamic social relationships. The negotiation pipeline opens the potential of communicating with others to prob the other’s mind or reach a consensus. The hierarchical strategy with reflection enhances the ability to handle long-term planning. The self-evolving mechanism (reflection with self-play memory) further improves the overall performance without manual supervision. These modules cover most of the challenges in multi-agent interactions. The potential applications of such an AI agent are vast, ranging from simulated diplomatic environments to real-world assistance and analysis. In future research, we intend to develop a more realistic game space, characterized by incomplete information and multi-player games, to enhance and refine our model further. We will also extend the framework to other multi-agent scenarios, including embodied interactions (Zhong et al., 2023; Ci et al., 2023; Chen et al., 2023), sensor networks (Wang et al., 2021; Xu et al., 2020; Pan et al., 2022; Li et al., 2020), and video games (Wang et al., 2024a; Ma et al., 2024). This framework can also be employed to develop various applications. For instance, in the fields of economics and finance, we intend to utilize it to create business analytics and negotiation models.
Code Release
Implementation of the proposed method can be found at this link for reproducing our results: https://github.com/todexter3/Richelieu.git
Acknowledgements
This work was supported by the National Science and Technology Major Project (MOST-2022ZD0114900), NSF 6240073415, and Qualcomm University Research Grant.
@@ A.1 Rules of Diplomacy Game
  • The number of units a country can have cannot exceed the number of supply centers it controls. If the number of supply centers decreases, excess units must be disbanded. Each autumn, new units can be built at supply centers. Coastal supply centers can produce fleets or armies, while others can only produce armies. Hill [2014]
@@ A.2 Domain Knowledge
Richelieu can adopt a strategy of allying with distant countries while attacking neighboring ones to occupy adjacent territories and achieve rapid expansion. Richelieu should pay attention to the Balance of Power by forming alliances with other countries or supporting weaker states to prevent any single country or alliance from becoming too powerful. David [2014] To this end, Richelieu can also adopt a strategy of attacking distant countries while allying with nearby ones, sacrificing short-term benefits to avoid the emergence of future hegemonic states that could threaten his own survival. When facing multiple enemies, Richelieu can find ways to divide other countries and incite wars among them. Whether in offense or defense, Richelieu should actively choose suitable allies. Richelieu can also introduce a third party to achieve goals such as ceasefire, alliance, or joint attack. To achieve alliances or ceasefires, Richelieu can sacrifice some interests to the other party as long as the ultimate benefits are greater. Others may lie and deceive Kostick [2015]; their words in negotiations are not binding. Richelieu must avoid being deceived or betrayed. At the same time, Richelieu can also actively deceive others to achieve his own goals.Richard [1979], Allan [1975]